# SelaVPR++: Towards Seamless Adaptation of Foundation Models for Efficient Place Recognition

### v1

@@ III-E Training Strategy and Unified VPR Training Dataset
In addition to SALAD-CM [25] (combining GSV-Cities and MSLS for training with a complex mining strategy), the latest work SuperPlace [72] also merges multiple datasets, i.e., GSV-Cities, SF-XL, MSLS, and Pitts250k (not Pitts30k as in ours), for model training and achieves excellent performance. Nevertheless, their processing methods for each dataset are different, hindering the integration of diverse datasets into a unified framework and even limiting the performance of trained models. For example, SuperPlace [72] and our work both use the CosPlace [28] method to assign place images into a finite number of categories to process datasets other than GSV-Cities. However, SuperPlace fails to fully utilize the angle information of the Pittsburgh (Pitts250k) and MSLS datasets. It requires to implement local feature matching for the Pittsburgh dataset to assist in category division. For MSLS, it compromises by assuming that all images have the same orientation, without using compass angles data. In summary, different from existing methods, we unify the framework of common training datasets with more consistent processing. This enables us to merge these datasets for better training of VPR models.
@@ IV-A Datasets and Performance Evaluation
Pitts30k is collected from Google Street View panoramas, and provides 24 images with different viewpoints at each place. So this dataset exhibits large viewpoint variations (and moderate condition changes). Pitts30k is a subset of Pitts250k (but more difficult than Pitts250k) and contains 10k database images each in the training, validation, and test sets.
Tokyo24/7 includes 75984 database images and 315 query images captured from urban scenes. The query images are selected from 1125 images taken at 125 distinct places with 3 different viewpoints and at 3 different times of day. This dataset mainly exhibits viewpoint changes and drastic condition changes (day-night changes).
Mapillary Street-Level Sequences (MSLS) is a large-scale VPR dataset containing over 1.6 million images labeled with GPS coordinates and compass angles, captured from 30 cities in urban, suburban, and natural scenes over seven years. It covers various challenging visual changes due to illumination, weather, season, viewpoint, as well as dynamic objects, and includes subsets of training, public validation (MSLS-val), and withheld test (MSLS-challenge). Following several related works [12, 13, 24, 23], the MSLS-val and MSLS-challenge sets are used to evaluate models.
Nordland primarily consists of suburban and natural place images, captured from the same viewpoint in the front of a train across four seasons, which allows the images to show severe condition (e.g., season and weather) changes but no viewpoint variations. Its ground truth is provided by the frame-level correspondence. Following previous works [2, 75], we use the winter images as queries and the summer images for reference, each containing 27592 images.
@@ IV-B Implementation Details
We use DINOv2 (both the base and large versions) as the foundation model and conduct experiments on NVIDIA GeForce RTX 3090 GPUs using PyTorch. The resolution of the input image is 224$\times$224 in training and 322$\times$322 in inference. For the model using DINOv2-base, the token dimension in the backbone is 768 and the number of input channels for the MultiConv module is 384, that is, the bottleneck ratio of adapters is 0.5. The 1×1 convolution before the 3×3 and 5×5 convolution reduces the number of channels to 24. The numbers of output channels for the three convolutional paths (1×1, 3×3, and 5×5) are 192, 96, and 96, respectively. For the model using DINOv2-large, the token dimension in the backbone is 1024, and the number of channels for each convolution layer in the MultiConv adapter is also 1024/768 times higher than that for DINOv2-base. We set the hyperparameters of the MS loss as in [16]. The coefficient $\lambda$ in Eq. (15) is set to 0.1. We train our models using the Adam optimizer with the initial learning rate set as 0.0004, halved every 3 epochs. A training batch consists of 120 places with 4 images each (i.e., 480 images). When the performance on MSLS-val does not have improvement for 12 epochs, the training is terminated. We also set the maximum epochs to 25. Additionally, for the DINOv2-base backbone, we equip each transformer block (12 blocks in total) with a MultiConv adapter. However, for the DINOv2-large backbone, we only equip the last 16 transformer encoder blocks with MultiConv adapters. This enables them to be trained on a single NVIDIA 3090 GPU with the above large batch size.
@@
TABLE II: Comparison to State-of-the-Art Methods on VPR Benchmark Datasets. The Best Is Highlighted in Bold and the Second Is Underlined. $\dagger$ CricaVPR Uses a Cross-Image Encoder to Correlate Multiple Images from the Same Place to Achieve Better Performance on Pitts30k. It Is Not Included in the Comparison with Others on Pitts30k.
@@
Method Dim Pitts30k-test Tokyo24/7 MSLS-val MSLS-challenge Nordland Average
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
Patch-NetVLAD [11] / 88.7 94.5 95.9 86.0 88.6 90.5 79.5 86.2 87.7 48.1 57.6 60.5 44.9 50.2 52.2 69.4 75.4 77.4
TransVPR [12] / 89.0 94.9 96.2 79.0 82.2 85.1 86.8 91.2 92.4 63.9 74.0 77.5 63.5 68.5 70.2 76.4 82.2 84.3
SelaVPR [13] / 92.8 96.8 97.7 94.0 96.8 97.5 90.8 96.4 97.2 73.5 87.5 90.6 87.3 93.8 95.6 87.7 94.3 95.7
NetVLAD [6] 32768 81.9 91.2 93.7 60.6 68.9 74.6 53.1 66.5 71.1 35.1 47.4 51.7 6.4 10.1 12.5 47.4 56.8 60.7
SFRS [47] 4096 89.4 94.7 95.9 81.0 88.3 92.4 69.2 80.3 83.1 41.6 52.0 56.3 16.1 23.9 28.4 59.5 67.8 71.2
CosPlace [28] 512 88.4 94.5 95.7 81.9 90.2 92.7 82.8 89.7 92.0 61.4 72.0 76.6 58.5 73.7 79.4 74.6 84.0 87.3
MixVPR [44] 4096 91.5 95.5 96.3 85.1 91.7 94.3 88.0 92.7 94.6 64.0 75.9 80.6 76.2 86.9 90.3 81.0 88.5 91.2
EigenPlaces [75] 2048 92.5 96.8 97.6 93.0 96.2 97.5 89.1 93.8 95.0 67.4 77.1 81.7 71.2 83.8 88.1 82.6 89.5 92.0
CricaVPR [23] 4096 94.9 97.3 98.2 93.0 97.5 98.1 90.0 95.4 96.4 69.0 82.1 85.7 90.7 96.3 97.6 87.5 93.7 95.2
SALAD [24] 8448 92.5 96.4 97.5 94.6 97.5 97.8 92.2 96.4 97.0 75.0 88.8 91.3 89.7 95.5 97.0 88.8 94.9 96.1
SALAD-CM [25] 8448 92.7 96.8 97.9 96.8 97.5 97.8 94.2 97.2 97.4 82.7 91.2 92.7 96.0 98.5 99.2 92.5 96.2 97.0
BoQ [26] 12288 93.7 97.1 97.9 98.1 98.1 98.7 93.8 96.8 97.0 79.0 90.3 92.0 90.6 96.0 97.5 91.0 95.7 96.6
SelaVPR++ (resource) 2048 93.3 96.6 97.6 97.5 98.7 99.0 94.5 96.9 97.3 81.0 91.9 93.7 94.6 97.8 98.3 92.2 96.4 97.2
SelaVPR++ (performance) 4096 94.4 97.5 98.1 98.1 98.7 99.4 94.5 98.0 98.2 84.0 93.7 94.4 97.2 99.0 99.4 93.6 97.4 97.9
@@ IV-C Comparison with State-of-the-Art Methods
1) Quantitative results. Table II shows the quantitative comparison results and our performance-focused SelaVPR++ achieves the best R@1/R@5/R@10 on all datasets. The methods using the foundation model DINOv2 as the backbone all achieve outstanding performance and surpass previous works (e.g., CosPlace, MixVPR, and EigenPlaces) on these datasets that exhibit real-world diverse challenges. This fully demonstrates that the foundation model can provide powerful feature representation with appropriate fine-tuning. Moreover, it is worth mentioning that, except for our SelaVPR++, no single method can consistently outperform others on all datasets. Specifically, BoQ and SALAD-CM, as two SOTA methods, exhibit their own distinct strengths. BoQ shows significant advantages over others on urban datasets, i.e., Pitts30k and Tokyo24/7. In contrast, SALAD-CM achieves better performance on the datasets covering suburban and natural scenes, i.e., MSLS and Nordland. However, our performance-focused SelaVPR++ gets the best results on all datasets, although the dimension of our global descriptors is less than half of theirs and we use more compact binary features to retrieve candidates. For example, although BoQ and SALAD-CM have achieved 93.7% R@1 on Pitts30k and 96.0% R@1 on Nordland respectively, our method can further improve R@1 to 94.4% on Pitts30k and 97.2% on Nordland. More importantly, our method achieves 93.7% R@5 on MSLS-challenge and ranks first on the official leaderboard, which is a large improvement over the previous best result (91.2% of SALAD-CM). Besides, our SelaVPR++ significantly outperforms its predecessor, our previous SelaVPR work, by 5.9% for the average R@1 of five datasets. We also provide the comparison between our SelaVPR++ and other SOTA methods on some supplementary VPR datasets (i.e., SPED [35], Eynsham [76], and SVOX [77], all provided by [2]) as shown in Table III, and SelaVPR++ still gets the best results. These results suggest that our method can produce highly robust global descriptors and excel in various challenging scenarios.
@@
In addition, our resource-focused SelaVPR++ is also comprehensively superior to the vast majority of methods (e.g., EigenPlaces, CricaVPR, and SALAD) by a considerable margin, and obviously outperforms BoQ for the average R@1/R@5/R@10. Although the average R@1 of our resource-focused SelaVPR++ is 0.3% lower than that of SALAD-CM, ours achieves the best average R@5 and R@10. More importantly, compared to SALAD-CM (8448-dim) and BoQ (12288-dim), our resource-focused SelaVPR++ employs low-dimensional (2048-dim) global descriptors for re-ranking and more compact (512-dim) binary descriptors for initial retrieval. This indicates that our resource-focused model can achieve significantly higher retrieval efficiency, which will be further demonstrated in subsequent ablation experiments.
@@
Moreover, to more intuitively illustrate that our SelaVPR++ can perform well in the VPR task, we visualize the attention map of the output feature map (before aggregation) from different models, including the pre-trained foundation model DINOv2, our previous SelaVPR work and the proposed SelaVPR++. As shown in Fig. 6, the pre-trained model tends to pay attention to dynamic foreground objects (e.g., pedestrians and vehicles), and often ignores some static discriminative backgrounds (e.g., buildings and vegetation) that are useful to identify places. SelaVPR ameliorates the disadvantage to some extent through global adaptation and focuses more on static buildings. However, SelaVPR still occasionally pays some attention to dynamic vehicles or leaves out useful buildings in some challenging cases. SelaVPR++ can perform better than SelaVPR in details and obviously distinguish static discriminative landmarks from the dynamic foreground, which is important for a robust VPR model. This demonstrates that SelaVPR++ can better bridge the gap between the model pre-training and VPR tasks compared to SelaVPR.
Figure 6: Attention map visualizations of the pre-trained foundation model (DINOv2), our previous work SelaVPR, and the proposed SelaVPR++. SelaVPR aims to seamlessly adapt pre-trained models to VPR. However, it does not do well in details for these examples. In the first five examples, it still focuses on some dynamic vehicles. Besides, in the last two examples, SelaVPR does not pay enough attention to the partial discriminative buildings in the right region of the image or even ignores them. In contrast, SelaVPR++ can significantly filter out the dynamic foreground and focus on static landmarks useful for VPR.

### v2

@@ III-E Training Strategy and Unified VPR Training Dataset
In addition to SALAD-CM [26] (combining GSV-Cities and MSLS for training with a complex mining strategy), the latest work SuperPlace [76] also merges multiple datasets, i.e., GSV-Cities, SF-XL, MSLS, and Pitts250k (not Pitts30k as in ours), for model training and achieves excellent performance. Nevertheless, their processing methods for each dataset are different, hindering the integration of diverse datasets into a unified framework and even limiting the performance of trained models. For example, SuperPlace [76] and our work both use the CosPlace [30] method to assign place images into a finite number of categories to process datasets other than GSV-Cities. However, SuperPlace fails to fully utilize the angle information of the Pittsburgh and MSLS datasets. It requires to implement local feature matching for the Pittsburgh dataset to assist in category division. For MSLS, it compromises by assuming that all images have the same orientation, without using compass angles data. In summary, different from existing methods, we unify the framework of common training datasets with more consistent processing. This enables us to merge these datasets for better training of VPR models.
@@ IV-A Datasets and Performance Evaluation
Pitts30k is collected from Google Street View panoramas, and provides 24 images with different viewpoints at each place. So this dataset shows large viewpoint changes. Pitts30k is a subset of Pitts250k (but harder than Pitts250k).
Tokyo24/7 are captured from urban scenes and contains about 76k images in database. This dataset mainly exhibits viewpoint changes and drastic (day-night) condition changes.
MSLS is a large-scale VPR dataset containing over 1.6 million images labeled with GPS coordinates and compass angles, captured from 30 cities in urban, suburban, and natural scenes over 7 years. It covers various challenging visual changes due to illumination, weather, season, viewpoint, as well as dynamic objects, and includes subsets of training, public validation (MSLS-val), and withheld test (MSLS-challenge). The MSLS-val and MSLS-challenge sets are used to evaluate models.
Nordland primarily consists of suburban and natural place images, captured from the same viewpoint in the front of a train across four seasons, which allows the images to show severe condition (e.g., season) changes but no viewpoint variations. Following previous works [2, 79], we use the winter images as queries and the summer images for reference.
@@ IV-B Implementation Details
We use DINOv2 (both the base and large versions) as the foundation model and conduct experiments on NVIDIA GeForce RTX 3090 GPUs using PyTorch. The resolution of the input image is $\times224$ in training and $\times322$ in inference. For the model using DINOv2-base, the token dimension in the backbone is 768 and the number of input channels for the MultiConv module is 384, that is, the bottleneck ratio of adapters is 0.5. The 1×1 convolution before the 3×3 and 5×5 convolution reduces the number of channels to 24. The numbers of output channels for the three convolutional paths (1×1, 3×3, and 5×5) are 192, 96, and 96, respectively. For the model using DINOv2-large, the token dimension in the backbone is 1024, and the number of channels for each convolution layer in the MultiConv adapter is also 1024/768 times higher than that for DINOv2-base. We set the hyperparameters of the MS loss as in [16]. The coefficient $\lambda$ in Eq. (15) is set to 0.1. We train models using the Adam optimizer with the initial learning rate set as 0.0004, halved every 3 epochs. A training batch consists of 120 places with 4 images each (i.e., 480 images). When the performance on MSLS-val does not have improvement for 12 epochs, the training is terminated. We also set the maximum epochs to 25. Additionally, for the DINOv2-base backbone, we equip each transformer block (12 blocks in total) with a MultiConv adapter. However, for the DINOv2-large backbone, we only equip the last 16 transformer encoder blocks with MultiConv adapters. This enables them to be trained on a single NVIDIA 3090 GPU with the above large batch size.
@@
TABLE II: Comparison to State-of-the-Art Methods on VPR Benchmark Datasets. The Best Is Highlighted in Bold and the Second Is Underlined. $\dagger$ CricaVPR Uses a Cross-Image Encoder to Correlate Multiple Images from the Same Place to Achieve Better Performance on Pitts30k. It Is Not Included in the Comparison with Others on Pitts30k. The Results of BoQ Are Measured by Us Using the Official Code and Model Weights.
@@
Method Dim Backbone Training Pitts30k-test Tokyo24/7 MSLS-val MSLS-challenge Nordland Avg.
Dataset R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@5
Patch-NetVLAD [11] / VGG16 / 88.7 94.5 95.9 86.0 88.6 90.5 79.5 86.2 87.7 48.1 57.6 60.5 44.9 50.2 52.2 75.4
TransVPR [12] / / / 89.0 94.9 96.2 79.0 82.2 85.1 86.8 91.2 92.4 63.9 74.0 77.5 63.5 68.5 70.2 82.2
SelaVPR [13] / DINOv2-L / 92.8 96.8 97.7 94.0 96.8 97.5 90.8 96.4 97.2 73.5 87.5 90.6 87.3 93.8 95.6 94.3
NetVLAD [6] 32768 VGG16 Pitts30k 81.9 91.2 93.7 60.6 68.9 74.6 53.1 66.5 71.1 35.1 47.4 51.7 6.4 10.1 12.5 56.8
SFRS [49] 4096 VGG16 Pitts30k 89.4 94.7 95.9 81.0 88.3 92.4 69.2 80.3 83.1 41.6 52.0 56.3 16.1 23.9 28.4 67.8
CosPlace [30] 512 VGG16 SF-XL 88.4 94.5 95.7 81.9 90.2 92.7 82.8 89.7 92.0 61.4 72.0 76.6 58.5 73.7 79.4 84.0
MixVPR [46] 4096 ResNet50 GSV 91.5 95.5 96.3 85.1 91.7 94.3 88.0 92.7 94.6 64.0 75.9 80.6 76.2 86.9 90.3 88.5
EigenPlaces [79] 2048 ResNet50 SF-XL 92.5 96.8 97.6 93.0 96.2 97.5 89.1 93.8 95.0 67.4 77.1 81.7 71.2 83.8 88.1 89.5
CricaVPR [24] 4096 DINOv2-B GSV 94.9 97.3 98.2 93.0 97.5 98.1 90.0 95.4 96.4 69.0 82.1 85.7 90.7 96.3 97.6 93.7
SALAD [25] 8448 DINOv2-B GSV 92.5 96.4 97.5 94.6 97.5 97.8 92.2 96.4 97.0 75.0 88.8 91.3 89.7 95.5 97.0 94.9
SALAD-CM [26] 8448 DINOv2-B GSV+MSLS 92.7 96.8 97.9 96.8 97.5 97.8 94.2 97.2 97.4 82.7 91.2 92.7 96.0 98.5 99.2 96.2
BoQ [80] 12288 DINOv2-B GSV 93.7 97.1 97.9 96.5 97.8 98.4 93.8 96.8 97.0 79.0 90.3 92.0 90.6 96.0 97.5 95.6
SelaVPR++(resource) 2048 DINOv2-B Unified 93.3 96.6 97.6 97.5 98.7 99.0 94.5 96.9 97.3 81.0 91.9 93.7 94.6 97.8 98.3 96.4
SelaVPR++(performance) 4096 DINOv2-L Unified 94.4 97.5 98.1 98.1 98.7 99.4 94.5 98.0 98.2 84.0 93.7 94.4 97.2 99.0 99.4 97.4
@@ IV-C Comparison with State-of-the-Art Methods
1) Quantitative results. Table II shows the quantitative comparison results and our performance-focused SelaVPR++ achieves the best R@1/R@5/R@10 on all datasets. The methods using the foundation model DINOv2 as the backbone all achieve outstanding performance and surpass previous works (e.g., CosPlace, MixVPR, and EigenPlaces) on these datasets that exhibit real-world diverse challenges. This fully demonstrates that the foundation model can provide powerful feature representation with appropriate fine-tuning. Moreover, it is worth mentioning that, except for our SelaVPR++, no single method can consistently outperform others on all datasets. Specifically, BoQ and SALAD-CM, as two SOTA methods, exhibit their own distinct strengths. BoQ shows significant advantages over others on urban datasets, i.e., Pitts30k and Tokyo24/7. In contrast, SALAD-CM achieves better performance on the datasets covering suburban and natural scenes, i.e., MSLS and Nordland. However, our performance-focused SelaVPR++ gets the best results on all datasets, although the dimension of our global descriptors is less than half of theirs and we use more compact binary features to retrieve candidates. For example, although BoQ and SALAD-CM have achieved 93.7% R@1 on Pitts30k and 96.0% R@1 on Nordland respectively, our method can further improve R@1 to 94.4% on Pitts30k and 97.2% on Nordland. More importantly, our method achieves 93.7% R@5 on MSLS-challenge and ranks first on the official leaderboard, which is a large improvement over the previous best result (91.2% of SALAD-CM). Besides, our SelaVPR++ significantly outperforms its predecessor, our previous SelaVPR work, by 3.1% for the average R@5 of five datasets. We also provide the comparison between our SelaVPR++ and other SOTA methods on some supplementary VPR datasets (i.e., SPED [37], Eynsham [81], and SVOX [82], all provided by [2]) as shown in Table III, and SelaVPR++ still gets the best results. These results suggest that our method can produce highly robust global descriptors and excel in various challenging scenarios.
@@
In addition, our resource-focused SelaVPR++ is also comprehensively superior to the vast majority of methods (e.g., EigenPlaces, CricaVPR, and SALAD) by a considerable margin, and outperforms BoQ and SALAD-CM for the average R@5. More importantly, compared to SALAD-CM (8448-dim) and BoQ (12288-dim), our resource-focused SelaVPR++ employs low-dimensional (2048-dim) global descriptors for re-ranking and more compact (512-dim) binary descriptors for initial retrieval. This indicates that our resource-focused model can achieve significantly higher retrieval efficiency, which will be further demonstrated in subsequent ablation experiments. It is worth mentioning that, even though our performance-focused SelaVPR++ uses a larger DINOv2-large backbone, its resource requirements are less than SALAD and BoQ (with the DINOv2-base backbone). We have measured the GPU memory usage of all three methods during training, and the results are summarized in Table IV. Our method consumes nearly half the memory compared to SALAD, and even less than BoQ, which only fine-tunes the last two blocks.
To ensure a fair comparison with BoQ and SALAD, as well as show the scalability of SelaVPR++, we additionally conduct experiments with more consistent settings, including the same backbone (DINOv2-base), training dataset, descriptor dimensionality, and image resolution ($\times224$ in training, different from $\times280$ in the BoQ paper [80]). The first way to unify the dimensionality is to reduce the descriptor dimensionalities of BoQ and SALAD to the same as our method (2048-dim). This can be easily achieved by reducing the dimensionality of the linear layers in BoQ and the number of clusters in SALAD. The results are shown in the first three rows of Table V. SelaVPR++ outperforms SALAD (2048-dim) and BoQ (2048-dim) on all datasets. The second way to unify the dimensionality is to increase the descriptor dimensionality of SelaVPR++ to the same as SALAD and BoQ, respectively. Given that the default GeM pooling in SelaVPR++ is not suitable for high dimensionalities, we correspondingly replace GeM in SelaVPR++ with the BoQ/SALAD aggregator for comparison. Since our method is not in conflict with BoQ and SALAD, this is also easy to achieve. The results are shown in the remaining rows of Table V, and we can summarize three points: (1) With the same training dataset, our method almost always outperforms BoQ/SALAD, which further highlights the strength of our method. (2) On the Pitts30k, MSLS-val, and Nordland datasets, both SelaVPR++(BoQ) and SelaVPR++(SALAD) trained on the unified dataset outperform SelaVPR++(resource) in Table II, showing that SelaVPR++ can be further improved with stronger aggregators, i.e., good scalability. (3) Other SOTA methods can also benefit a lot from our unified dataset, indicating that it is a universal contribution to VPR.
TABLE IV: The Training GPU Memory Usage of Our Method, BoQ, and SALAD. We Consistently Set the Training Image Size to $224\times 224$ and the Batch Size to 40 for All Methods.
[…]
SelaVPR++ DINOv2-L Our adaptation method 7.97
BoQ DINOv2-B Partial-tuning(2 blocks) 8.20
SALAD DINOv2-B Partial-tuning(4 blocks) 14.91
TABLE V: Consistent Comparison to SALAD/BoQ. All Methods Use DINOv2-Base Backbone. SALAD (2048-Dim) and BoQ (2048-Dim) Are the Dimensionality-Reduced Versions of SALAD and BoQ, Which Are the Same 2048-Dim as Our SelaVPR++. SelaVPR++(SALAD) and SelaVPR++(BoQ) Use the SALAD and BoQ Aggregators to Replace GeM Pooling to Output Descriptors of the Same Dimensionality as SALAD and BoQ, Respectively, i.e., 8448-Dim and 12288-Dim. The Image Resolution Is $224\times 224$ in Training and $322\times 322$ in Inference.
Method Training Pitts30k-test MSLS-val Tokyo24/7 Nordland
Dataset R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
SALAD (2048-dim) GSV 91.5 96.1 91.2 96.2 93.0 96.5 77.2 87.8
BoQ (2048-dim) 92.0 96.2 91.8 96.1 94.9 97.5 77.7 88.7
SelaVPR++ 92.6 96.5 92.6 96.2 96.2 97.8 81.3 91.3
SALAD GSV 92.5 96.4 92.2 96.4 94.6 97.5 89.7 95.5
SelaVPR++(SALAD) 93.2 96.9 93.5 97.0 95.6 97.8 89.9 95.5
BoQ 92.5 96.5 92.2 96.4 96.2 98.4 84.8 93.3
SelaVPR++(BoQ) 93.3 96.8 92.4 96.4 96.2 98.1 92.1 96.6
SALAD Unified 93.0 96.6 94.5 97.2 95.6 97.8 94.6 98.2
SelaVPR++(SALAD) 93.8 97.3 94.6 97.6 96.5 98.1 96.7 98.9
BoQ 93.0 96.5 94.8 97.4 96.2 98.4 96.5 98.8
SelaVPR++(BoQ) 94.1 97.1 94.9 97.6 96.8 98.4 97.6 99.2
Figure 6: Heatmap visualizations of feature maps produced by the pre-trained foundation model (DINOv2), our previous work SelaVPR, and the proposed SelaVPR++. We compute the mean in the channel dimension of the output feature map and display it using the heatmap. SelaVPR aims to seamlessly adapt pre-trained models to VPR. However, it does not do well in detail for these examples. In the first two examples, it is still disturbed by a few dynamic vehicles. Besides, in the last two examples, SelaVPR does not pay enough attention to the partial discriminative buildings in the right region of the image or even ignores them. In contrast, SelaVPR++ can significantly filter out the dynamic foreground and focus on static landmarks useful for VPR.
@@
Moreover, to more intuitively illustrate that our SelaVPR++ can perform well in the VPR task, we visualize the attention map of the output feature map (before aggregation) from different models, including the pre-trained foundation model DINOv2, our previous SelaVPR work and the proposed SelaVPR++. As shown in Fig. 6, the pre-trained model tends to pay attention to dynamic foreground objects (e.g., pedestrians and vehicles), and often ignores some static discriminative backgrounds (e.g., buildings and vegetation) that are useful to identify places. SelaVPR ameliorates the disadvantage to some extent through global adaptation and focuses more on static buildings. However, SelaVPR still occasionally pays some attention to dynamic vehicles or leaves out useful buildings in some challenging cases. SelaVPR++ can perform better than SelaVPR in detail and obviously distinguish static discriminative landmarks from the dynamic foreground, which is important for a robust VPR model. This demonstrates that SelaVPR++ can better bridge the gap between the model pre-training and VPR tasks compared to SelaVPR.